# PaddleOCR-VL: Boosting Multilingual Document Parsing via a 0.9B Ultra-Compact Vision-Language Model

### v2

@@ 1 Introduction
The inherent complexity of modern documents presents unique challenges: they often combine dense text, complex tables or chart, mathematical expressions, multiple languages and handwritten texts, with deserve layout structures. Recent research [1, 9, 10, 11, 12] in the field of document parsing primarily following two technological approaches. The first approach [9, 10] employs pipeline methodologies based on specialized, modular expert models. Although these methods offer strong performance, they are increasingly hindered by integration complexity, cumulative error propagation, and inherent limitations when handling highly complex documents. Secondly, end-to-end approaches [12, 13, 14] leveraging multimodal models aim to simplify the workflow and enable joint optimization. However, these methods often struggle with correct text order and can even generate hallucinations when faced with lengthy or complex layouts, while also incurring substantial computational overhead for long sequence outputs, thereby restricting their practical deployment.
@@ 4.1 Page-level Evaluation
Table 2 demonstrate that PaddleOCR-VL achieves SOTA performance, outperforming existing pipeline tools, general VLMs, and other specialized document parsing models across all key metrics. Specifically, our model achieves a top-ranking overall score of 92.56, surpassing the next best model, MinerU2.5-1.2B (90.67). Moreover, our model establishes new SOTA results in the sub-tasks, including the lowest Text-Edit distance [44] of 0.035, the highest Formula-CDM score of 91.43, the leading scores of 89.76 and 93.52 in Table-TEDS and Table-TEDS-S, and the best readering ordering scores of 0.043, respectively. These results underscore its superior accuracy in text recognition, formula recognition, and complex table structure analysis.
@@ 4.3 Inference Performance
Table 13 provides a comprehensive comparison of inference efficiency across different methods. The proposed PaddleOCR-VL demonstrates clear and consistent advantages in both processing speed and memory efficiency. When deployed with the vLLM backend, it achieves 15.8% higher page throughput and 14.2% higher token throughput than the leading baseline, MinerU2.5, establishing itself as the most efficient solution overall. In addition, PaddleOCR-VL achieves notable memory savings, using roughly 40% less GPU memory than dots.ocr while sustaining significantly faster processing. These results collectively confirm that PaddleOCR-VL attains state-of-the-art inference efficiency through a balanced optimization of speed and memory usage, making it highly suitable for real-world, high-throughput document understanding scenarios.
@@
MonkeyOCR-pro-1.2B [1] 1456.4 0.6730 1120.3 75.5
dots.ocr [52] 2784.6 0.3522 532.9 78.5
MinerU2.5 [2] 927.3 1.0574 1647.9 41.9
PaddleOCR-VL 800.9 1.2241 1881.2 43.7
PaddleOCR-VL 917.6 1.0684 1641.5 49.8
Table 13: End-to-End Inference Performance Comparison. denotes the vLLM backend, and denotes the SGLang backend.
@@ A.4 Chart
    Public Data Collection and Cleaning: We collected a large number of samples from public datasets, including ChartQA [32], PlotQA [33], Chart2Text [34], DVQA [35], Unichart [36], Beagle [37], ChartINFO [38], visText [39], and ExcelChart [40]. However, the raw datasets suffered from poor annotation quality and extremely imbalanced data distributions. Thus, a meticulous data cleaning and filtering pipeline was implemented to remove noisy samples and ensure balanced clustering, resulting in a high-quality dataset of 220k samples.
@@ Appendix C Inference Performance on Different Hardware Configurations
We measured the inference performance of PaddleOCR-VL on different hardware configurations, as summarized in Table A2. As observed, PaddleOCR-VL demonstrates stable and efficient inference performance across a wide range of hardware and backend configurations, showing that the system can flexibly adapt to diverse computing environments. Moreover, we are currently integrating the FastDeploy backend, which is expected to further enhance inference efficiency in future releases.
@@
A100 vLLM 800.9 1.2241 1881.2 43.7
SGLang 917.6 1.0684 1641.5 49.8
A10 vLLM 1238.0 0.7921 1217.2 14.1
SGLang 1429.9 0.6858 1055.8 20.0
RTX 3060 vLLM 2749.1 0.3568 548.2 11.9
SGLang 2792.4 0.3513 540.8 11.8
RTX 5070 vLLM 1292.9 0.7584 1165.5 8.9
RTX 4090D vLLM 845.3 1.1597 1781.8 16.7
SGLang 951.8 1.0303 1586.1 21.8

### v3

@@ 1 Introduction
The inherent complexity of modern documents presents unique challenges: they often combine dense text, complex tables or chart, mathematical expressions, multiple languages and handwritten texts, with diverse layout structures. Recent research [1, 9, 10, 11, 12] in the field of document parsing primarily following two technological approaches. The first approach [9, 10] employs pipeline methodologies based on specialized, modular expert models. Although these methods offer strong performance, they are increasingly hindered by integration complexity, cumulative error propagation, and inherent limitations when handling highly complex documents. Secondly, end-to-end approaches [12, 13, 14] leveraging multimodal models aim to simplify the workflow and enable joint optimization. However, these methods often struggle with correct text order and can even generate hallucinations when faced with lengthy or complex layouts, while also incurring substantial computational overhead for long sequence outputs, thereby restricting their practical deployment.
@@ 4.1 Page-level Evaluation
Table 2 demonstrate that PaddleOCR-VL achieves SOTA performance, outperforming existing pipeline tools, general VLMs, and other specialized document parsing models across all key metrics. Specifically, our model achieves a top-ranking overall score of 92.86, surpassing the next best model, MinerU2.5-1.2B (90.67). Moreover, our model establishes new SOTA results in the sub-tasks, including the lowest Text-Edit distance [44] of 0.035, the highest Formula-CDM score of 91.22, the leading scores of 90.89 and 94.76 in Table-TEDS and Table-TEDS-S, and the best readering ordering scores of 0.043, respectively. These results underscore its superior accuracy in text recognition, formula recognition, and complex table structure analysis.
@@ 4.3 Inference Performance
Table 13 provides a comprehensive comparison of inference efficiency across different methods. The proposed PaddleOCR-VL demonstrates clear and consistent advantages in processing speed. When deployed with the FastDeploy [^1] https://github.com/PaddlePaddle/FastDeploy backend, it achieves 53.1% higher page throughput and 50.9% higher token throughput than the leading baseline, MinerU2.5, establishing itself as the most efficient solution overall. These results collectively confirm that PaddleOCR-VL attains state-of-the-art inference efficiency through a balanced optimization of speed, making it highly suitable for real-world, high-throughput document understanding scenarios.
@@
MonkeyOCR-pro-1.2B [1] vLLM (v0.10.2) 1456.4 0.6730 1120.3
dots.ocr [52] vLLM (v0.10.2) 2784.6 0.3522 532.9
MinerU2.5 [2] vLLM (v0.10.2) 927.3 1.0574 1647.9
PaddleOCR-VL SGLang (v0.5.2) 882.1 1.1115 1707.8
PaddleOCR-VL vLLM (v0.10.2) 728.7 1.3453 2067.6
PaddleOCR-VL FastDeploy (v2.3) 605.6 1.6184 2486.4
Table 13: End-to-End Inference Performance Comparison.
@@ A.4 Chart
    Public Data Collection and Cleaning: We collected a large number of samples from public datasets, including ChartQA [32], PlotQA [33], Chart2Text [34], DVQA [35], Unichart [36], Beagle [37], ChartINFO [38], visText [39], and ExcelChart [40]. However, the raw datasets suffered from poor annotation quality and extremely imbalanced data distributions. Inspired by [76], a meticulous data cleaning and filtering pipeline was implemented to remove noisy samples and ensure balanced clustering, resulting in a high-quality dataset of 220k samples.
@@ Appendix C Inference Performance on Different Hardware Configurations
We measured the inference performance of PaddleOCR-VL on different hardware configurations, as summarized in Table A2. As observed, PaddleOCR-VL demonstrates stable and efficient inference performance across a wide range of hardware and backend configurations, showing that the system can flexibly adapt to diverse computing environments.
@@
H800 FastDeploy 400.4 2.2250 3416.7 63.5
vLLM 596.7 1.6442 2524.9 44.2
SGLang 875.2 1.1198 1720.8 49.5
A100 FastDeploy 605.6 1.6184 2486.4 62.8
vLLM 728.7 1.3453 2067.6 40.1
SGLang 882.1 1.1115 1707.8 49.7
H20 FastDeploy 635.0 1.5432 2371.2 64.3
vLLM 694.7 1.4112 2168.8 74.8
SGLang 759.6 1.2906 1986.6 80.2
L20 FastDeploy 802.0 1.2224 1881.9 40.8
vLLM 904.9 1.0835 1665.1 25.2
SGLang 953.6 1.0280 1579.0 31.3
A10 FastDeploy 1197.2 0.8191 1260.9 21.8
vLLM 1320.5 0.7426 1141.2 15.6
SGLang 1446.2 0.6781 1043.8 19.9
RTX 3060 vLLM 2694.1 0.3641 559.5 11.9
SGLang 2765.2 0.3547 546.1 11.9
RTX 4090D vLLM 851.9 1.1507 1768.1 16.6
SGLang 932.4 1.0514 1618.5 20.9